INAF queue observing, 2016 Mar 02 - Mar 03

Observer: R. Carini, M. Faccini
Telescope Operator: David Gonzalo-Huerta
Support Astronomer: O. Kuhn on call
Instrument: LUCI, LBC, PARALLEL

Summary

at the sunset the sky is almost clear with some thin clouds

Night duration (12\xB0 el twilight) 19:44-5:27 ( 02:44-12:27 UT) = 10 h 00m

Observing time: 10 h (100%)

Weather time loss: 0 h ()

technical time loss 0 m ()

Data Summary

Project_RUN Objects Filter ExpTime Notes
Raimondo Uma II   1h 16 m  
Castellano MACSJ0717 U_spec; 1h 30m last ob: 10
Bellazzini DDT Secco C2 dual grating; r_sloan 1h completed
Bellazzini DDT Secco C1 dual grating; r_sloan 1 h COMPLETED PROGRAM
standard Hz44 dual grating    
Antoniucci PVcep dual grating 20m to be repeated

Details

00:55 sky is almost clear with some thin clouds
  • start dark calibration for LUCI2
  • good luci2.20160303.0004
  • 01:19 the TO has to open the dome to thermalize.

01:48 start LBC flat field.
  • LBC stucked

02:25 Raimondo program
  • run copointing UMaII _f10_UBr_point.ob
  • dofpia
  • lbcrangebal *dofpia

02:42 start science. run UMaII _f10_UBr.ob
  • the first image in the red (lbcr.20160303.024925) presents a ghost in the chip3 ( left chip), maybe caused by a lamp turned on at the LBTO at level 4
  • 3:03 seeing ~0.8'''

03:32 we take the missing images of the field 8 for the program of Raimondo (with texp=30 sec)

3:38 dofpia on UMaII _f11_UBr.ob
  • run science UMaII _f11_UBr.ob *3:51 seeing ~0.7''-0.9'''

04:32 track error restart CMU

04:35 slew to Castellano program, target A
  • dofpia
  • lbcrangebal
  • 04:50 run science Castellano_MACSJ0717_UV_8.ob
  • 04:56 seeing ~0.9''
  • 05:03 seeing ~0.9''

05:29 dofpia on Castellano_MACSJ0717_UV_9.ob
  • 05:36 run science
  • 05:42 seeing ~0.8''
  • 06:06 seeing ~0.8'''
  • 6:07 the image lbcb.20160303,055339 presents dark lateral bands in every chip

06:20 slew to ob 10 for the program of Castellano
  • LBC stucked. the TO restart LBC

06:26 dofpia on Castellano_MACSJ0717_UV_10.ob
  • 6:30 run Castellano_MACSJ0717_UV_10.ob
  • 06:45 seeing ~0.7''-0.8''

07:11 Photometric standard S101

07:28 switch to Parallel mode: LBC-red + MODS1 *07:55 strt with the target C2
  • acqMODS SECCO_C2.acq mods1r.20160303.0001-0003
  • dofpia
  • 08:17 execMODS SECCO_C2.obs mods1r.20160303.0004-0006,mods1b.20160303.0001-0003
  • run ob of LBC SECCO_C2_1.ob
  • 08:24 seeing ~0.8''-0.9''
  • dofpia
  • 08:47 run SECCO_C2_2.ob
  • 09:16 dofpia
  • 09:20 run SECCO_C2_3.ob

09:45 slew to the target C1
  • acqMODS SECCO_C1.acq mods1r.20160303.0007-0009
  • dofpia
  • 10:10 execMODS SECCO_C1.obs mods1r.20160303.0010-0012,mods1b.20160303.0004-0006
  • run SECCO_C1_1.ob
  • 10:17 seeing ~1.0''
  • 10:38 dofpia
  • 10:42 SECCO_C1_2.ob, seeing ~1.0''
  • 11:0 airmass 1.50
  • 11:07 SECCO_C1_3.ob, seeing ~1.0''

11:29 slew to standard Hz44 *acqMods Hz44.acq mods1r.20160303.0013
  • 11:37 problems with the collimation

11:39 acqMODS HZ44.acq mods1r.20160303.0014-0015 *11:55 execMODs HZ44.acq.obs mods1r.20160303.0016-0018mods,1b.20160303.0007-0009

12:05 slew to Antoniucci program, target PVcep
  • acqMODS pvcep_ls.acq mods1r.20160303.0019-0022
  • execMODS pvcep_ls.obs mods1r.20160303.0023-0026mods,1b.20160303.0010-0013
  • the images in the blue grating are not good

1:02 LBC flat field
  • problems with LBC

*1:30 LUCI Dark
  • calibration with MODS
  • flat slitless dual grating
  • flat with slit: 0.8''(Antoniucci), 1.2''(Bellazzini), 5''(standard)

-- %USERSIG{RobertaCarini - 2016-03-03}%

Comments

 
Topic revision: r6 - 10 Nov 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback